21. Consider a university with a population mean GPA of 2.95, a standard deviation of 2, and a sampling distribution of all possible samples of size 100. The mean of the sample means is equal to 2.95 e) 025
Answer #1

Answer

Population mean GPA is equal to 2.95

Standard deviation of GPA is 0.20

we know that the standard deviation changes with the sample size, but the mean remains same for population as well as for sample mean.

So, for a sample size of 100

Mean GPA= 2.95

and standard deviation = o/V n = 0.2/v/100 0.2/10 0.02

So, correct answer for mean of sample means is 2.95

Option D is correct